Suppose Qd=-50*P+806, Qs=37*P+18 for home and Qd*=-50*P+136, Qs*=37*P+30 for foreign. What is the free trade price to two decimal places?
Answer #1

Answer:

Given:

For Home: Qd=-50*P+806, Qs=37*P+18

For Foreign : Qd*=-50*P+136, Qs*=37*P+30

Total quantity demanded (Q'): Qd + Qd*

Q' = -50P + 806 – 50P + 136 = -100P + 942

Total quantity supplied (Q''): Qs + Qs*

Q'' = 37P + 18 + 37P + 30 = 74P + 48

Equilibrium is attained when Q' = Q"

-100P + 942 = 74P + 48

174P = 894

P = 5.137 (Free trade price)

Thus free trade price is 5.14